OpenStudy (anonymous):

What is the vertex of the graph y=x^2+5x+6

OpenStudy (anonymous):

the vertex of the graph y=x^2+5x+6 a=1,b=5,c=6 vertex x =-b/2a=-5/2(1)=-5/2 subs this to y=x^2+5x+6 y=x^2+5x+6 y=(-5/2)^2+5(-5/2)+6=-1/4 vertex=(-5/2, -1/4) ans